OpenStudy (anonymous):

Give a formula for the extended function that is continuous at the indicated point. f(x)=(x-4)/(sqrt(x)-2), x=4

OpenStudy (mathmath333):

\(\Large\rm \lim_{x\to 4}\dfrac{x-4}{\sqrt x-2}=f(4)\) \(\Large\rm =\lim_{x\to 4}\dfrac{(\sqrt x)^2-2^2}{\sqrt x-2}\) \(\Large\rm =\lim_{x\to 4}\dfrac{[(\sqrt x)-2][\sqrt x)+2]}{\sqrt x-2}\) \(\Large\rm =\lim_{x\to 4}\sqrt x+2\) \(\Large\rm =\sqrt 4+2\) \(\Large\rm =2+2=4\) \(\Large\rm or\) \(\Large\rm =-2+2=0\) \(\Large\rm f(x)=\begin{cases} \dfrac{x-4}{\sqrt x-2}, x\ne 4\\\\ x=4,~~~~ x=4\\ x=0,~~~~x=4 \end{cases}\)